Graph the system below and write its solution
3x+y=-6
y=-1/2x-1

Answer by mananth(16946)

y=-3x-6
y= -3 x -6
Generate values for y from values of x
x 2 4 -4
y= -12 -18 6
(x,y) (2,0) (4,6) ,(-4,-18)


y= - 1/ 2 x -1
Generate values for y from values of x
x 2 4 -4
y= -2 -3 1
(x,y) (2,-2) (4,-3) ,(-4,1)





(-2,0)